OpenStudy (amonoconnor):

I am looking to have my answer verified in regard to the integral of the following: [((1 - sin(2t))^(3/2))(cos(2t))]dt, from x = 0 to (pi/4). My answer: 1/10, or .1 Any and all insight, or help, is greatly appreciated! :)

hartnn (hartnn):

how did u get 1/10 ? i am getting 1/5

hartnn (hartnn):

made a u substitution u = 1-sin 2t right ?

OpenStudy (amonoconnor):

... Yep, you're right. In simplifying the anti-derivative, I was careless, and did some faulty mental math. 2/10 = 1/5, not 1/10. My bad, and thank you very much! I assume that by changing that, the rest of the algebra changes proportionally, and will result in a final answer of 1/5, or .2

hartnn (hartnn):

thats correct :)

OpenStudy (amonoconnor):

And yes, that is the method I used:)

OpenStudy (amonoconnor):

Thank you very much! :)

hartnn (hartnn):

welcome ^_^